# Songrun He   Linying Lv   Asaf Manela   Jimmy Wu

### v2

@@ 4 Conclusion
In this paper, we address the critical challenge of lookahead and survivorship bias in LLMs used in social science applications, particularly in financial forecasting. By introducing ChronoBERT and ChronoGPT, two series of chronologically consistent language models trained on timestamped text data, we demonstrate that chronological consistency can be achieved without compromising performance. Our results show that ChronoBERT and ChronoGPT match or surpass the language comprehension abilities of BERT and GPT-2 while generating performance comparable to the much larger Llama 3.1 model in asset pricing applications.
@@
We make our ChronoBERT and ChronoGPT models publicly available at: https://huggingface.co/manelalab.
Our work suggests several potential avenues for future research. One direction would be developing compute-optimal training strategies specifically tailored for chronologically consistent LLMs. While we have demonstrated that strong language understanding can be achieved without introducing lookahead bias, further work is needed to establish scaling laws (akin to Chinchilla scaling laws from Hoffmann et al. (2022)) that account for the unique constraints of temporal data limitations (Muennighoff et al., 2023). Such scaling laws would provide guidance on the optimal allocation of computational resources when training models with historical data cutoffs, potentially revealing different optimal ratios of parameters to training tokens compared to models trained on all available data.
@@ A.2 Hellaswag Evaluation
Appendix B Architectural Differences Between BERT and GPT
@@
B.1 Pretraining Objectives
B.1.1 BERT: Masked Language Modeling (MLM)
@@
B.1.2 GPT: Causal Language Modeling (CLM)
@@ B.2 Attention Mechanisms
B.2.1 BERT: Bidirectional Self-Attention
@@
B.2.2 GPT: Causal Self-Attention
@@
B.3 Sequence Classification Approaches
B.3.1 BERT for Sequence Classification

### v3

@@ 4 Conclusion
In this paper, we address the critical challenge of lookahead and training leakage in LLMs used in social science applications, particularly in financial forecasting. By introducing ChronoBERT and ChronoGPT, two series of chronologically consistent language models trained on timestamped text data, we demonstrate that chronological consistency can be achieved without compromising performance. Our results show that ChronoBERT and ChronoGPT match or surpass the language comprehension abilities of BERT and GPT-2 while generating performance comparable to the much larger Llama 3.1 model in asset pricing applications.
@@
Our work suggests several potential avenues for future research. One avenue would be to assess lookahead bias in other settings. For this purpose, we make our ChronoBERT and ChronoGPT models publicly available at: https://huggingface.co/manelalab. Another avenue would be developing compute-optimal training strategies specifically tailored for chronologically consistent LLMs. While we have demonstrated that strong language understanding can be achieved without introducing lookahead bias, further work is needed to establish scaling laws (akin to Chinchilla scaling laws from Hoffmann et al. (2022)) that account for the unique constraints of temporal data limitations (Muennighoff et al., 2023). Such scaling laws would provide guidance on the optimal allocation of computational resources when training models with historical data cutoffs, potentially revealing different optimal ratios of parameters to training tokens compared to models trained on all available data.
@@ A.2 Hellaswag Evaluation
Appendix B Scaling Law of ChronoGPT
A critical question is whether performance can be improved despite the inherent limitations of the chronological training data. According to scaling laws (Kaplan et al., 2020), model performance is a function of parameter count, data size, and compute budget. In our context, the latter two factors face a bottleneck. Scaling the training dataset is challenging due to the strict requirement of chronological consistency. Furthermore, with this constrained dataset, training beyond a certain point (e.g., four epochs) yields modest benefit on loss reduction.
Motivated by the success of model scaling demonstrated by OpenAI (Radford et al., 2019) and AllenAI (OLMo et al., 2024), we investigate if increasing model size is an effective strategy within our data-constrained regime. We test four architectures: 124M, 353M, 1B, and 1.5B parameters, detailed in Table B.1.
Model Size Vocabulary Layers Heads $d_{\text{model}}$
124 M 50,304 12 6 768
353 M 50,304 24 8 1,024
1 B 50,304 36 12 1,536
1.5 B 50,304 52 12 1,536
Table B.1: Architecture hyperparameters for the four ChronoGPT model sizes.
This table summarizes the core architectural hyperparameters of each scaled-up ChronoGPT model.
We then evaluate them on validation loss and language understanding tasks against our baseline, using a consistent knowledge cutoff of December 1999. As shown in Figure B.1, larger models consistently achieve lower validation loss and higher HellaSwag scores, demonstrating that scaling up model size improves both general language modeling and specific language understanding capabilities. The 1.5B parameter model achieves the best performance across both metrics, with validation loss converging to approximately 1.32 after 80 billion training tokens, compared to 1.35 for the smallest 124M model. Similarly, the 1.5B model reaches a HellaSwag score of 0.37, substantially outperforming the 353M model (0.32) and the smallest 124M model (0.295).
Figure B.1: Validation Loss and Language Understanding Performance Comparison
This figure compares the cross-entropy loss and HellaSwag scores of ChronoGPT models with four sizes: 124 million (124M), 353 million (353M) parameters, 1 billion (1B) parameters, and 1.5 billion (1.5B) parameters. Validation losses are reported in logs. All models have a consistent knowledge cutoff of December 1999.
These results confirm that scaling model size is an effective strategy for improving both language modeling and downstream task performance, even within a data-constrained regime. Based on these prominent gains and considering our compute budget, we select the 1.5B parameter architecture to train the full ChronoGPT series, extending its knowledge base from 1999 to 2024.
Appendix C Architectural Differences Between BERT and GPT
@@
C.1 Pretraining Objectives
C.1.1 BERT: Masked Language Modeling (MLM)
@@
C.1.2 GPT: Causal Language Modeling (CLM)
@@ C.2 Attention Mechanisms
C.2.1 BERT: Bidirectional Self-Attention
@@
C.2.2 GPT: Causal Self-Attention
@@
C.3 Sequence Classification Approaches
C.3.1 BERT for Sequence Classification